# Optimizing Age of Information in Networks with Large and Small Updates

### v1

@@ III Lower Bound
Waiting Time and Service Time. Let $K_{i}(T)$ denote the total number of delivered updates from source $i$ by the end of slot $T$ and let $m\in\{1,...,K_{i}(T)\}$ be the index of the delivered updates from source $i$. Let $t^{\prime}_{i}[m]$ denote the time slot in which the first packet of the $m$th delivered update is received, and let $t_{i}[m]$ denote the time slot in which the last packet of the $m$th delivered update is received. We define the waiting time of the $m$th update from source $i$ as $W_{i}[m]\;:=\;t^{\prime}_{i}[m]-t_{i}[m-1]$, which is the interval between the delivery of the last packet of the $(m-1)$th update and the delivery of the first packet of the $m$th update. Similarly, the service time of the $m$th update for source $i$ is defined as $S_{i}[m]\;:=\;t_{i}[m]-t^{\prime}_{i}[m]$, which is the interval between the delivery of the first and last packets of the $m$th update. We assume that $t_{i}[0]=0$, $W_{i}[0]=0$, and $S_{i}[0]=0$ for all $i$.
@@
[…]
@@ V Max-Weight Policy
For a network with parameters $\{N,\alpha_{i},p_{i},L_{i}\}$, by choosing the constants $\beta_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}},\gamma_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}\sqrt{p_{i}}}$ and $\bar{q}_{i}=q^{L_{B}}_{i}-\epsilon$, where $\epsilon\rightarrow 0$, the optimality ratio of Max-Weight policy is such that
$\rho=6+\frac{\sqrt{\varPsi}}{NL_{B}}$ (33)
@@
$-\frac{L_{i}\sqrt{p_{i}}}{\bar{q}_{i}}+2\frac{L^{2}_{i}}{\bar{q}^{2}_{i}}+2\frac{L_{i}^{2}}{\bar{q}_{i}}\Bigl{]}\Biggl{)}$
@@ VI Simulation Results
[…]
Figure 7: Simulation results for networks with a varying number of sources. The network comprises $N$ sources with $N\in\{10,20,30,50,70,100\}$. For each source, the weight is drawn from $\alpha_{i}\sim U[1,10]$, the channel reliability from $p_{i}\sim U[0.5,1]$, and each source is equally likely to belong to Class 1 (with $L_{i}\sim U[2,5]$) or Class 2 (with $L_{i}\sim U[20,100]$).
@@
[…]
@@ Appendix A Proof of Propsition 1
The evolution of $h_{i}(t)$ is well-defined in each of the time intervals $W_{i}[m]$, $S_{i}[m]$, and $R_{i}$ According to (3), during the interval$\big{[}t_{i}[m-1]+1,t_{i}[m]\big{]}$, the parameter $h_{i}(t)$ evolves as $\{S_{i}[m-1]+2,S_{i}[m-1]+3,...,S_{i}[m-1]+W_{i}[m]+S_{i}[m]+1\}$. This pattern is repeated throughout the entire time-horizon, for $m\in\{1,2,...,K_{i}(T)\}$, and also during the last $R_{i}$ slots. As a result, the time-average AoI associated with destination $i$ can be expressed as
@@ Appendix F Upper Bound for Lyapunov Drift
$z_{i}(t+1)+L_{i}(t+1)=\begin{cases}z_{i}(t)+L_{i}(t),&\text{if }d_{i}(t)=1\text{ and }L_{i}(t)>1\\ &\text{ or }L_{i}(t)=L,\\ L_{i}+1,&\text{if }d_{i}(t)=1\text{ and }L_{i}(t)=1,\\ z_{i}(t)+L_{i}(t)+1,&\text{otherwise.}\end{cases}$ (103)
@@ Appendix G Proof of Theorem 9
Theorem 9. For any given network model with network parameters $\{N,\alpha_{i},p_{i},L_{i}\}$, by choosing the constant $\beta_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}},\gamma_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}\sqrt{p_{i}}}$ and $\bar{q}_{i}=q^{L_{B}}_{i}-\epsilon$, where $\epsilon\rightarrow 0$, the optimality ratio of Max-Weight policy is such that
$\rho=6+\frac{\sqrt{\varPsi}}{NL_{B}}$ (104)
where
$\varPsi=\left(\sum_{i=1}^{N}\alpha_{i}\right)\left(\sum_{i=1}^{N}\frac{\alpha_{i}}{\sqrt{p_{i}}}\left[5\frac{L^{2}_{i}\sqrt{p_{i}}}{\bar{q}^{2}_{i}}-\frac{L_{i}}{\bar{q}_{i}}-\frac{L_{i}\sqrt{p_{i}}}{\bar{q}_{i}}+2\frac{L^{2}_{i}}{\bar{q}^{2}_{i}}+2\frac{L_{i}^{2}}{\bar{q}_{i}}\right]\right)$ (105)
@@
$B(t)=$ $\sum_{i=1}^{N}\beta_{i}\mathbb{I}_{L_{i}(t)=L}\left[2h_{i}(t)-1\right]$
$+$ $\sum_{i=1}^{N}\gamma_{i}\mathbb{I}_{L_{i}(t)>1}\left[2\left(z_{i}(t)+L_{i}(t)\right)-1\right]+V\left[x^{+}_{i}(t)\bar{q}_{i}+\frac{1}{2}\right],$
@@
$+\gamma_{i}\mathbb{I}_{L_{i}(t)>1}\left[2z_{i}(t)+2L_{i}(t)-1\right]$
Before we prove the upper bound of EWSAoI achieved by the Max-Weight policy, we first prove the lower bound for achievable throughput in the following Lemma 11.
@@
The Max-Weight policy achieves any feasible set of minimum throughput targets $\{\bar{q}\}^{N}_{i=1}$, if there exists any SRP $s$ satisfies $p_{i}\mu^{s}_{i}\geq\bar{q},\forall i$.
@@
Summarizing (28) and take the average over the time-horizon yields
$\frac{\mathbb{E}[\mathcal{L}(T+1)]}{T}-\frac{\mathbb{E}[\mathcal{L}(1)]}{T}\leq\frac{1}{T}\sum_{t=1}^{T}B(t)-\frac{1}{TN}\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i}\mathbb{E}\left[u_{i}(t)|\mathbb{S}(t)\right]C_{i}(t),$ (106)
and taking the limit as $T\to\infty$, we obtain
$\frac{1}{N}\sum_{i=1}^{N}p_{i}\mathbb{E}\left[u_{i}(t)C_{i}(t)\right]\leq\mathbb{E}\left[B(t)\right]$ (107)
Recall that the Max-Weight policy minimizes the RHS of (29) by selecting $i=arg\max{C_{i}(t)}$ in every slot $t$. Hence, any other policy $\pi\in\Pi$ yields a lower (or equal) RHS. Consider a SRP $s\in\Pi^{s}_{r}$ that, in each slot $t$, selects node $i$ with probability $\mu^{s}_{i}\in(0,1]$. The scheduling decision of policy $s$ is independent with network state, yields
where $u^{S}_{i}(t)$ denote the scheduling decision made by optimal SRP $S$.
The furhter analysis yields
$\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i}\mathbb{E}\left[u_{i}(t)|\mathbb{S}(t)\right]C_{i}(t)\geq\frac{1}{T}\sum_{t=1}^{T}\sum_{i=1}^{N}p_{i}\mathbb{E}\left[u^{S}_{i}(t)|\mathbb{S}(t)\right]C_{i}(t)$ (109)
where both sides of (109) are the extended expectation, which can be simplified as
$\sum_{i=1}^{N}p_{i}\mathbb{E}\left[u_{i}(t)C_{i}(t)\right]\geq\sum_{i=1}^{N}p_{i}\mu_{i}^{s}\mathbb{E}\left[C_{i}(t)\right]$ (110)
Substitute (30) and (31) into (110) and rearrange the terms, we obtain (111) on the top of next page. For simplicity of exposition, we divide inequality (111) into six terms $LHS_{1}+LHS_{2}+LHS_{3}\leq RHS_{1}+RHS_{2}+RHS_{3}$, gives
$\begin{gathered}\sum_{i=1}^{N}p_{i}\mu^{s}_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=1}]\mathbb{E}\left[\beta_{i}\left(h^{2}_{i}(t)-2h_{i}(t)z_{i}(t)\right)+\gamma_{i}\left(\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}\right)\mid L_{i}(t)=1\right]+\sum_{i=1}^{N}V\left(p_{i}\mu^{s}_{i}-q_{i}\right)x_{i}^{+}(t)\\ \leq\sum_{i=1}^{N}\beta_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=L}]\mathbb{E}\left[2h_{i}(t)-1\mid L_{i}(t)=L\right]+\sum_{i=1}^{N}\gamma_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)>1}]\mathbb{E}\left[2z_{i}(t)+2L_{i}(t)-1\mid L_{i}(t)>1\right]+\frac{NV}{2}\end{gathered}$ (111)
$LHS_{1}=\sum_{i=1}^{N}p_{i}\mu^{s}_{i}\beta_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=1}]\mathbb{E}\left[h^{2}_{i}(t)-2h_{i}(t)z_{i}(t)\mid L_{i}(t)=1\right]$ (112)
$LHS_{2}=\sum_{i=1}^{N}p_{i}\mu^{s}_{i}\gamma_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=1}]\mathbb{E}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}\mid L_{i}(t)=1\right]$ (113)
$LHS_{3}=\sum_{i=1}^{N}V\left(p_{i}\mu^{s}_{i}-q_{i}\right)\mathbb{E}[x_{i}^{+}(t)]$ (114)
$RHS_{1}=\sum_{i=1}^{N}\beta_{i}(1-p_{i}\mu_{i})\mathbb{E}[\mathbb{I}_{L_{i}(t)=L}]\mathbb{E}\left[2h_{i}(t)-1\mid L_{i}(t)=L\right]$ (115)
$RHS_{2}=\sum_{i=1}^{N}\gamma_{i}(1-p_{i}\mu_{i})\mathbb{E}[\mathbb{I}_{L_{i}(t)>1}]\mathbb{E}\left[2z_{i}(t)+2L_{i}(t)-1\mid L_{i}(t)>1\right]$ (116)
$RHS_{3}=\frac{NV}{2}$ (117)
Since that $LHS_{1}>-\infty$, $LHS_{2},RHS_{1}>,RHS_{2},RHS_{3}$ are positive, we establish that
$\lim_{T\rightarrow}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[x_{i}^{+}(t)]<\infty,$ (118)
which is sufficient to establish that the throughput achieved by Max-Weight policy satisfies that $q_{i}^{MW}\geq\bar{q}_{i},\forall i$[26, Theorem 2.8]
Next, we leverage Lemma 11 to prove the upper bound for EWSAoI associate with Max-Weight policy in Theorem 9.
Consider a set of target throughput subject to
@@
where $\sigma\rightarrow 0$. By setting $\{\bar{q}_{i}\}_{i=1}^{N}$ as the target value in the throughput debt $x^{+}_{i}(t)$, since there exists a SRP $\mu_{i}^{L_{B}}=q_{i}^{L_{B}}/p_{i}$, the throughput achieve by Max-Weight policy, i.e. $\{q^{MW}_{i}\}_{i=1}^{N}$, satisfies
@@
Substitute (31) into (107), notice that $q^{MW}_{i}=p_{i}\mathbb{E}\left[u_{i}(t)\right]$ we obtain $LHS^{\prime}_{1}+LHS^{\prime}_{2}+LHS^{\prime}_{3}\leq RHS^{\prime}_{1}+RHS^{\prime}_{2}+RHS^{\prime}_{3}$, which are given by
$LHS^{\prime}_{1}=$ $\sum_{i=1}^{N}q^{MW}_{i}\beta_{i}\mathbb{E}\left[\mathbb{I}_{L_{i}(t)=1}|u_{i}(t)=1\right]$ (121)
×$\times\mathbb{E}\left[h^{2}_{i}(t)-2h_{i}(t)z_{i}(t)|u_{i}(t)=1,L_{i}(t)=1\right]$
$LHS^{\prime}_{2}=$ $\sum_{i=1}^{N}q^{MW}_{i}\gamma_{i}\mathbb{E}\left[\mathbb{I}_{L_{i}(t)=1}|u_{i}(t)=1\right]$ (122)
×$\times\mathbb{E}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}|u_{i}(t)=1,L_{i}(t)=1\right]$
$LHS^{\prime}_{3}=\sum_{i=1}^{N}V\left(q^{MW}_{i}-\bar{q}_{i}\right)\mathbb{E}[x_{i}^{+}(t)]$ (123)
$RHS^{\prime}_{1}=\sum_{i=1}^{N}\beta_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=L}]\mathbb{E}\left[2h_{i}(t)-1\mid L_{i}(t)=L\right]$ (124)
$RHS^{\prime}_{2}=\sum_{i=1}^{N}\gamma_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)>1}]\mathbb{E}\left[2z_{i}(t)+2L_{i}(t)-1\mid L_{i}(t)>1\right]$ (125)
@@
Next, we analysis the packet transmission process and the bound for each term to simplify (121) – (126).
Notice that the sufficient and necessary condition for a update delivery of source $i$ is that $d_{i}(t)\mathbb{I}_{L_{i}(t)=1}=1$, and the time average update delivery of source $i$ is $\frac{q_{i}}{L_{i}}$, then we establish
$\frac{q_{i}}{L_{i}}=\mathbb{E}[d_{i}(t)\mathbb{I}_{L_{i}(t)=1}]$ (127)
Expending the expectation on the RHS yields
@@
Notice that $q_{i}=p_{i}\mathbb{P}\left[u_{i}(t)=1\right]$, further substituting (127) into (128) yields
@@
Recall that Max-Weight policy selects the source with the highest value of $C_{i}(t)$ at each given time slot. Thus we establish
@@
$\mathbb{E}\left[C_{i}(t)|u_{i}(t)=1,L_{i}(t)=1\right]\geq\mathbb{E}\left[C_{i}(t)|L_{i}(t)=1\right]$ (131)
Due to the sequential packets transmission, during each update delivery process, the value $h_{i}(t)-z_{i}(t)$ and $z_{i}(t)+L_{i}(t)$ is nondecreasing until the delivery of last packet, regardless of the channel and scheduling decision, so we establish
$\mathbb{E}\left[C_{i}(t)|u_{i}(t)=1,L_{i}(t)=1\right]\geq\mathbb{E}\left[C_{i}(t)|u_{i}(t)=1\right]$ (132)
Similarly, we obtain
$\mathbb{E}\left[2h_{i}(t)-1\mid L_{i}(t)=L\right]\leq\mathbb{E}\left[2h_{i}(t)-1\right]$ (133)
$\mathbb{E}\left[2z_{i}(t)+2L_{i}(t)-1\mid L_{i}(t)>1\right]\leq\mathbb{E}\left[2z_{i}(t)+2L_{i}(t)-1\right]$ (134)
Substitute (129), (132) and (130) into (121) and (122), we obtain
@@
$\begin{array}[]{ll}LHS^{\prime}_{1}+LHS^{\prime}_{2}&\stackrel{{\scriptstyle(a)}}{{\geq}}\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\mathbb{E}\left[h^{2}_{i}(t)-2h_{i}(t)z_{i}(t)\right]+\sum_{i=1}^{N}\frac{q_{i}^{MW}\gamma_{i}}{L_{i}}\mathbb{E}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}|L_{i}(t)=1\right]\\ &\stackrel{{\scriptstyle(b)}}{{\geq}}\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\mathbb{E}\left[\left(h_{i}(t)-2z_{i}(t)\right)^{2}\right]+\sum_{i=1}^{N}\frac{q_{i}^{MW}\gamma_{i}}{L_{i}}\mathbb{E}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}|L_{i}(t)=1\right]\\ &\stackrel{{\scriptstyle(c)}}{{\geq}}\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\left[\mathbb{E}\left[h_{i}(t)\right]-2\mathbb{E}\left[z_{i}(t)\right]\right]^{2}+\sum_{i=1}^{N}\frac{q_{i}^{MW}\gamma_{i}}{L_{i}}\left(\mathbb{E}\left[z_{i}(t)+2|L_{i}(t)=1\right]^{2}-\left(L_{i}+1\right)^{2}\right)\\ &\stackrel{{\scriptstyle(d)}}{{\geq}}\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\left[\mathbb{E}\left[h_{i}(t)\right]-2\mathbb{E}\left[z_{i}(t)\right]\right]^{2}\end{array}$ (135)
where $(a)$ from (130), (131) and (132); $(b)$ from $h_{i}(t)>h_{i}(t)-2z_{i}(t)$; $(c)$ by applying Jensen Inequality $\mathbb{E}\left[\left(\bar{h}_{i}(t)-2z_{i}(t)\right)^{2}\right]\geq\left[\mathbb{E}\left[h_{i}(t)\right]-2\mathbb{E}\left[z_{i}(t)\right]\right]^{2}$, where $\bar{h}_{i}(t)=\max(h_{i}(t),2z_{i}(t))$ and $\mathbb{E}\left[\bar{h}_{i}(t)\right]\geq\mathbb{E}\left[h_{i}(t)\right]$; and $(d)$ from
$\mathbb{E}\left[z_{i}(t)+2|L_{i}(t)=1\right]\geq L_{i}+1$ (136)
Notice that $LHS^{\prime}_{2}+LHS^{\prime}_{3}>0,\mathbb{E}[\mathbb{I}_{L_{i}(t)=L}]<1,\mathbb{E}[\mathbb{I}_{L_{i}(t)>1}]<1$, substituting (133), (134) and (135) into $LHS^{\prime}_{1}+LHS^{\prime}_{2}+LHS^{\prime}_{3}\leq RHS^{\prime}_{1}+RHS^{\prime}_{2}+RHS^{\prime}_{3}$ yields
$\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\left[\mathbb{E}\left[h_{i}(t)\right]-2\mathbb{E}\left[z_{i}(t)\right]\right]^{2}$ (137)

### v2

@@ III Lower Bound
Waiting Time and Service Time. Let $K_{i}(T)$ denote the total number of delivered updates from source $i$ by the end of slot $T$ and let $m\in\{1,...,K_{i}(T)\}$ be the index of the delivered updates from source $i$. Let $t^{\prime}_{i}[m]$ denote the time slot in which the first packet of the $m$th delivered update is received, and let $t_{i}[m]$ denote the time slot in which the last packet of the $m$th delivered update is received. We define the waiting time of the $m$th update from source $i$ as $W_{i}[m]\;:=\;t^{\prime}_{i}[m]-t_{i}[m-1]$, which is the interval between the delivery of the last packet of the (m1)(m-1)th update and the delivery of the first packet of the $m$th update. Similarly, the service time of the $m$th update for source $i$ is defined as $S_{i}[m]\;:=\;t_{i}[m]-t^{\prime}_{i}[m]$, which is the interval between the delivery of the first and last packets of the $m$th update. We assume that $t_{i}[0]=0$, $W_{i}[0]=0$, and $S_{i}[0]=0$ for all $i$.
@@
$\quad+\;\frac{\bar{\mathbb{M}}\bigl[S_{i}[m-1]\,(W_{i}[m]+S_{i}[m])\bigr]}{\bar{\mathbb{M}}\bigl[W_{i}[m]+S_{i}[m]\bigr]}\;+1\Biggr),$
@@ V Max-Weight Policy
For a network with parameters $\{N,\alpha_{i},p_{i},L_{i}\}$, by choosing the constants $\beta_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}},\gamma_{i}=\frac{\alpha_{i}}{q_{i}^{L_{B}}\sqrt{p_{i}}}$, $V>0$ small enough (see Lemma 12), and $\bar{q}_{i}=q^{L_{B}}_{i}-\epsilon$, where $\epsilon\rightarrow 0$, the optimality ratio of Max-Weight policy is such that
$\rho^{MW}=6+\frac{\sqrt{\varPsi}}{NL_{B}}$ (33)
@@
$-\frac{L_{i}\sqrt{p_{i}}}{\bar{q}_{i}}+2\frac{L^{2}_{i}}{\bar{q}^{2}_{i}}+2\frac{L_{i}^{2}}{\bar{q}_{i}}+\frac{-L_{i}^{2}+8L_{i}+24}{2}\Bigl]\Biggl)$
@@ VI Simulation Results
Figure 4: Simulation results for networks with varying channel reliabilities. The network comprises $N=10$ sources equally divided into Class 1 with $\alpha_{i}=5$ and $L_{i}=2$ and Class 2 with $\alpha_{i}=1$ and $L_{i}=50$. Channel reliabilities vary over $p_{i}\in\{0.2,\,0.15,\,\ldots,\,1\}$ for all sources.
Figure 5: Simulation results for networks with varying update lengths. The network comprises $N=10$ sources equally divided into Class 1 with $\alpha_{i}=5$, $p_{i}=0.8$, and $L_{i}=2$ and Class 2 with $\alpha_{i}=1$ and $p_{i}=0.4$. Update lengths for Class 2 are parameterized as $L^{*}-2,\,L^{*}-1,\,L^{*},\,L^{*}+1,\,L^{*}+2]$ with $L^{*}\in\{15,\,20,\,\ldots,\,100\}$.
Figure 6: Simulation results for networks with varying weights. The network comprises $N=10$ sources equally divided into Class 1 with variable priorities $\alpha_{i}\in\{2,\,4,\,\ldots,\,20\}$ with fixed $L=2$ and $p_{i}=0.8$, and Class 2 sources have fixed parameters $\alpha_{i}=1$, $L_{i}=50$, and $p_{i}=0.4$.
Figure 7: Simulation results for networks with a varying number of sources. The network comprises NN sources with $N\in\{10,20,30,50,70,100\}$. For each source, the weight is drawn from $\alpha_{i}\sim U[1,10]$, the channel reliability from $p_{i}\sim U[0.5,1]$, and each source is equally likely to belong to Class 1 (with $L_{i}\sim U[2,5]$) or Class 2 (with $L_{i}\sim U[20,100]$).
@@
In Figures 4, 5, and 6, we consider networks with $N=10$ sources, equally divided into five Class 1 and five Class 2 sources. In Fig. 4, Class 1 sources are configured with a priority $\alpha_{i}=5$ and a fixed update length $L_{i}=2$, while Class 2 sources have $\alpha_{i}=1$ and $L_{i}=50$; in this setup, the channel reliabilities for all sources vary over the set $p_{i}\in\{0.2,0.15,\ldots,1\}$. In Fig. 5, the parameters for Class 1 sources remain the same ($\alpha_{i}=5$, $p_{i}=0.8$, $L_{i}=2$), and for Class 2 sources, while the priority is fixed at $\alpha_{i}=1$ and the channel reliability at $p_{i}=0.4$, the update lengths vary as $L^{*}-2,\,L^{*}-1,\,L^{*},\,L^{*}+1,\,L^{*}+2]$ with $L^{*}$ taking values from $\{15,\,20,\,\ldots,\,100\}$. In Fig. 6, Class 1 sources have variable priorities $\alpha_{i}^{*}\in\{2,\,4,\,\ldots,\,20\}$ with $p_{i}=0.8$ and $L_{i}=2$, while Class 2 sources have $\alpha_{i}=1$, $p_{i}=0.4$, and $L_{i}=50$.
@@ Appendix A Proof of Propsition 1
[…]
@@ Appendix F Upper Bound for Lyapunov Drift
[…]
@@ Appendix G Proof of Theorem 9
$B(t)=$ $\sum_{i=1}^{N}\beta_{i}\mathbb{I}_{L_{i}(t)=L}\left[2h_{i}(t)-1\right]+V\left[x^{+}_{i}(t)\bar{q}_{i}+\frac{1}{2}\right]+\sum_{i=1}^{N}\gamma_{i}\mathbb{I}_{1<L_{i}(t)<L_{i}}\left[2\left(z_{i}(t)+L_{i}(t)\right)-1\right],$
@@
$+\gamma_{i}\mathbb{I}_{1<L_{i}(t)<L_{i}}\left[2z_{i}(t)+2L_{i}(t)-1\right]+\gamma_{i}\mathbb{I}_{L_{i}(t)=1}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}\right]+Vx^{+}(t).$
Step 1: A throughput lemma for Max-Weight. Prior to deriving the upper bound on the EWSAoI achieved by the Max-Weight policy, we prove Lemma 11.
@@
If there exists an SRP $s$ with long-term packet throughput greater than the throughput target for all sources $i$, i.e., $p_{i}\mu^{s}_{i}\geq\bar{q_{i}},\forall i$, then the long-term packet throughput of the Max-Weight policy is also such that $q_{i}^{MW}\geq\bar{q_{i}},\forall i$.
@@
We compare MW with an arbitrary stationary randomized policy (SRP) $s$ and show that MW yields a no-larger drift term. We then substitute this bound into the drift expression, take expectations, sum over time, and let the horizon $T\to\infty$ to prove that the throughput debt process remains finite. By applying the stability result in [26, Theorem 2.8], we conclude that the throughput targets are met, which establishes Lemma 11. The proof relies on Lyapunov-drift techniques similar to those in [13, Appendix B], and the full derivation is provided in Appendix H. ∎
Next, we leverage Lemma 11 to derive the upper bound on the EWSAoI associated with the Max-Weight policy described in Theorem 9.
Consider the throughput targets below
@@
with arbitrarily small $\sigma>0$. Clearly, $\{\bar{q}_{i}\}_{i=1}^{N}$ is feasible since a SRP with $\mu_{i}^{L_{B}}=q_{i}^{L_{B}}/p_{i}$ can meet the throughput target. By setting $\{\bar{q}_{i}\}_{i=1}^{N}$ as the target value in the throughput debt $x^{+}_{i}(t)$, leveraging Lemma 11, the throughput achieved by Max-Weight policy, i.e. $\{q^{MW}_{i}\}_{i=1}^{N}$, satisfies
@@
[…]
$LHS^{\prime}_{1}=$ $\lim_{T\rightarrow\infty}\frac{1}{TN}\sum_{t=1}^{T}\sum_{i=1}^{N}q^{MW}_{i}\beta_{i}\mathbb{E}\left[\mathbb{I}_{L_{i}(t)=1}|u_{i}(t)=1\right]\mathbb{E}\left[h^{2}_{i}(t)-2h_{i}(t)z_{i}(t)|u_{i}(t)=1,L_{i}(t)=1\right]$ (106)
$LHS^{\prime}_{2}=$ $\lim_{T\rightarrow\infty}\frac{1}{TN}\sum_{t=1}^{T}\sum_{i=1}^{N}q^{MW}_{i}\gamma_{i}\mathbb{E}\left[\mathbb{I}_{L_{i}(t)=1}|u_{i}(t)=1\right]\mathbb{E}\left[\left(z_{i}(t)+2\right)^{2}-\left(L_{i}+1\right)^{2}|u_{i}(t)=1\!,L_{i}(t)=1\!\right]$ (107)
$LHS^{\prime}_{3}=$ $\lim_{T\rightarrow\infty}\frac{1}{TN}\sum_{t=1}^{T}\sum_{i=1}^{N}V\left(q^{MW}_{i}-\bar{q}_{i}\right)\mathbb{E}[x_{i}^{+}(t)|u_{i}(t)=1,\!L_{i}(t)=1]$ (108)
$RHS^{\prime}_{1}=$ $\lim_{T\rightarrow\infty}\frac{1}{TN}\sum_{t=1}^{T}\sum_{i=1}^{N}\beta_{i}\mathbb{E}[\mathbb{I}_{L_{i}(t)=L}\left(2h_{i}(t)-1\right)]$ (109)
$RHS^{\prime}_{2}\!=$ $\!\lim_{T\!\rightarrow\infty}\!\frac{1}{TN}\!\sum_{t=1}^{T}\!\sum_{i=1}^{N}\!\gamma_{i}\mathbb{E}[\mathbb{I}_{1<L_{i}(t)<L_{i}}\!\left(2z_{i}(t)\!+\!2L_{i}(t)\!-\!1\right)]$ (110)
@@
Next, we analyze the packet transmission process and develop bounds for each term (106)–(111). Notice that a necessary and sufficient condition for each update delivery of source $i$ is given by the indicator $d_{i}(t)\mathbb{I}_{L_{i}(t)=1}=1$, and that the time average update delivery rate of source $i$ is $q_{i}/L_{i}$. Then, assuming ergodicity of the packet delivery process, we can establish
$\frac{q_{i}}{L_{i}}=\mathbb{E}[d_{i}(t)\mathbb{I}_{L_{i}(t)=1}]\;,$ (112)
expanding the expectation on the RHS yields
@@
Recall that $q_{i}=p_{i}\mathbb{P}\left[u_{i}(t)=1\right]$, further substituting (112) into (113) yields
@@
Since the Max-Weight policy selects the source with the highest value of $C_{i}(t)$ at any given network state $\mathbb{S}(t)$, we establish that
@@
Step 3: A refined ordering property for $C_{i}(t)$. Next we show that, for a low enough value of $V$, the value of $C_{i}(t)$ in (31) when the last packet of an update is successfully transmitted is no smaller than the value of $C_{i}(t)$ when earlier packets of the same update were successfully transmitted.
Lemma 12.
For a given source $i$, a given update index $m$, and for a sufficiently small $V>0$, the value of $C_{i}(t)$ in (31) when the last packet of update $m$ is successfully transmitted is no smaller than the value of $C_{i}(t)$ when earlier packets of update $m$ were successfully transmitted, i.e.,
$C_{i}\bigl(t_{i}[m]\bigr)\;\geq\;C_{i}(t),\forall t^{\prime}[m]\leq t\leq t_{i}[m],\text{ s.t. }u_{i}(t)c_{i}(t)=1.$ (116)
Proof.
We compare the value $C_{i}(t[m])$ with that at the first successful packet $C_{i}(t^{\prime}[m])$, as well as with its values at the intermediate packet-delivery slots. By performing a simple case split, bounding the resulting differences through algebraic manipulation, and choosing $V>0$ sufficiently small, we show that the last successful packet maximizes $C_{i}(t)$ over all successful packets of update $m$, thereby obtaining (116). Detailed derivations are provided in Appendix I. ∎
Step 4: An inequality on the conditional expectation of $C_{i}(t)$. We now use Lemma 12 to derive an inequality for the conditional expectations of $C_{i}(t)$, which will be instrumental for the subsequent drift analysis.
Lemma 13.
For each source $i$, the conditional expected value of $C_{i}(t)$ at the slots where the last packet of an update is delivered (i.e., $L_{i}(t)=1$ and $u_{i}(t)=1$) is no smaller than the conditional expected value of $C_{i}(t)$ over all slots in which source $i$ is scheduled (i.e., $u_{i}(t)=1$), namely
$\mathbb{E}\bigl[C_{i}(t)\mid L_{i}(t)=1,u_{i}(t)=1\bigr]\;\geq\;\mathbb{E}\bigl[C_{i}(t)\mid u_{i}(t)=1\bigr].$ (117)
Proof.
Using Lemma 12 and comparing the value of $C_{i}(t[m])$ with the average of $C_{i}(t)$ over all packet-delivery slots of the same update yields a per-update inequality. Summing this inequality over all updates, normalizing over time, and letting $T\to\infty$, assuming the ergodicity [^1] Under the Max–Weight scheduling policy, the joint network state $\mathbb{S}(t)\triangleq\{h_{i}(t),\,z_{i}(t),\,L_{i}(t)\}_{i=1}^{N}$ forms a discrete-time Markov chain that is irreducible and aperiodic. If the Max-Weight policy stabilizes the network, then $h_{i}(t)$, $z_{i}(t)$, and $L_{i}(t)$ remain bounded almost surely, thus the chain is positive recurrent. Irreducibility, aperiodicity, and positive recurrence together ensure the existence of a unique stationary distribution, and hence $\mathbb{S}(t)$ is ergodic. of $\mathbb{S}(t)$ turns time averages into steady-state expectations. Finally, using the relation $L_{i}\,\mathbb{E}[\mathbb{I}_{L_{i}(t)=1}\mathbb{I}_{u_{i}(t)=1}]=\mathbb{E}[\mathbb{I}_{u_{i}(t)=1}]$, we obtain (117). Detailed derivations are given in Appendix J. ∎
Step 5: Bounding the drift and concluding the EWSAoI bound. Substitute (114), (115), and  (117) into (106) and (107), we obtain
$LHS^{\prime}_{1}+LHS^{\prime}_{2}+LHS^{\prime}_{3}$ (118)
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}\lim_{T\to\infty}\!\frac{1}{T}\!\sum_{t=1}^{T}\sum_{i=1}^{N}\frac{q^{MW}_{i}\beta_{i}}{L_{i}}\,\mathbb{E}\!\left[\,h_{i}^{2}(t)\!-\!2h_{i}(t)z_{i}(t)\right]\!+\!\!\lim_{T\to\infty}\!\frac{1}{T}\!\sum_{t=1}^{T}\sum_{i=1}^{N}\frac{q^{MW}_{i}\gamma_{i}}{L_{i}}\!\mathbb{E}\!\left[(z_{i}(t)+2)^{2}\!\!-\!(L_{i}+1)^{2}\!\right]$